OpenStudy (ranya99):

An airplane, initially moving at 3.0 m/s along a runway, begins to accelerate at 3.6 m/s2. How far down the runway will it be in 20. seconds?

OpenStudy (anonymous):

how long hav u been waiting?

OpenStudy (anonymous):

just call somebody when nobody sees u....

OpenStudy (anonymous):

by just clicking at @ button....

OpenStudy (anonymous):

(v-u)/t =a u= 3 m/s a = 3.6 m/s^2 t=2 s

OpenStudy (anonymous):

if u put in the above relation u will get the answer..............

ganeshie8 (ganeshie8):

you may use the formula : \[s= ut + \dfrac{1}{2}at^2\]